# LocAgent: Graph-Guided LLM Agents for Code Localization

### v1

@@ 1 Introduction
Additionally, to facilitate a comprehensive evaluation of code localization methods, we introduce Loc-Bench, a new benchmark specifically designed for this task. Existing benchmarks like SWE-Bench present significant limitations: (1) they risk contamination through data overlap with LLM training sets Mündler et al. (2024), and (2) they primarily focus on bug fixing, lacking diversity in maintenance scenarios such as feature requests, performance optimizations, and security fixes. In contrast, Loc-Bench covers diverse scenarios and mitigates potential contamination concerns by incorporating more recent examples from popular Python repositories collected after known LLM training cutoff dates. We collected 660 examples from popular Python repositories, balanced across bug reports (282), feature requests (203), security issues (31), and performance problems (144), creating a diverse evaluation environment that better reflects real-world code maintenance challenges. Additionally, we provide tooling to continuously update the benchmark with new examples, allowing researchers to maintain a fresh evaluation dataset as models evolve and training data cutoffs advance.
@@ 3.2 Agent-guided Code Search
3.3 Open Source Model Fine-tuning
@@ 4.2 Dataset construction
| Loc-Bench | Bug Report | 282 |
| Feature Request | 203 | |
| (Totoal = 660) | Security Issue | 31 |
| Performance Issue | 144 |
@@ 5.5 Efficiency Analysis
Table 4 presents an efficiency analysis comparing agent-based methods in terms of cost and the number of agent interactions required. MoatlessTools demonstrates good cost-efficiency and requires relatively fewer rounds of interaction. However, the dense embeddings it uses make it difficult and slow to adapt to fast-evolving codebases. SWE-agent and Openhands also show moderate costs but still do not match the efficiency of LocAgent. For LocAgent with Claude-3.5, although more rounds of interaction are required, the cost remains lower than that of Openhands, illustrating the token efficiency of our tool’s outputs. LocAgent with fine-tuned Qwen models stands out for its superior efficiency. Qwen2.5-7B(ft) is the most cost-efficient option, requiring only $0.05 per example, while Qwen2.5-32B(ft) offers a more cost-effective alternative to Claude-3.5. These results highlight the potential of fine-tuned open-source models as efficient alternatives, providing an optimal balance of cost-effectiveness and performance that surpasses other methods.
@@ 5.6 Ablation Study
5.7 Evaluation Result on Loc-Bench
@@
| IR-Based | CodeRankEmbed | 70.72 | 77.81 | 58.86 | 63.02 | 38.52 | 41.76 |
| Agentless | Claude-3.5 | 67.33 | 67.33 | 52.39 | 52.39 | 42.68 | 42.68 |
| OpenHands | Claude-3.5 | 80.89 | 81.20 | 69.18 | 69.34 | 60.09 | 60.25 |
| SWE-agent | Claude-3.5 | 79.35 | 79.35 | 63.33 | 63.48 | 52.23 | 52.23 |
| LocAgent (Ours) | Qwen2.5-7B(ft) | 79.20 | 80.43 | 63.02 | 63.02 | 51.93 | 52.39 |
| Claude-3.5 | 84.59 | 87.06 | 72.42 | 72.57 | 60.25 | 62.10 | |
To ensure the robustness and the generalization of our methods and fine-tuned Qwen models and eliminate potential data leakage, we conduct an evaluation of this new dataset. Since the Loc-Bench contains examples that edit more than 1 file and less than 5 files, we assess file localization at top-5 and top-10 ranks and function/module localization at top-10 and top-15 ranks. Table 6 demonstrates that our fine-tuned Qwen2.5-7B model exhibits strong generalization capabilities, maintaining competitive performance compared to SWE-agent using more expensive and strong model Claude-3.5. These results highlight the practicality of the fine-tuned Qwen2.5-7B model for real-world applications. Despite being an open-source alternative, it achieves a performance comparable to Claude-3.5, supporting its feasibility as a cost-effective substitute for commercial models in practical scenarios.
Additionally, we evaluate the performance across four different difficulty categories. Figure 5 clearly shows that our method outperforms other methods in almost all categories of code localization. However, it also highlights a noticeable decrease in performance across the other three categories (Feature Request, Performance, and Security) compared to the Bug Report category. This performance gap likely reflects our training data distribution, which contained more bug report examples, potentially leading to scaffolds better optimized for bug localization tasks. This trend suggests that while our method is highly effective for bug report localization, there is still room for improvement in handling the other categories through more balanced training data and category-specific optimization strategies.
@@ B.1 Dataset construction details
B.2 Dataset statistics
The distribution of modified files, modules, and functions is shown in Figure 8.

### v2

@@ 1 Introduction
Additionally, to facilitate a comprehensive evaluation of code localization methods, we introduce Loc-Bench, a new benchmark specifically designed for this task. Existing benchmarks like SWE-Bench present significant limitations: (1) they risk contamination through data overlap with LLM training sets Mündler et al. (2024), and (2) they primarily focus on bug fixing, lacking diversity in maintenance scenarios such as feature requests, performance optimizations, and security fixes. In contrast, Loc-Bench covers diverse scenarios and mitigates potential contamination concerns by incorporating more recent examples from popular Python repositories collected after known LLM training cutoff dates. Additionally, we provide tooling to continuously update the benchmark with new examples, allowing researchers to maintain a fresh evaluation dataset as models evolve and training data cutoffs advance.
@@ 2.2 LLM-based Generative Retrieval Methods
2.3 Graph-based Code Representation Methods
| Method | Relation Types | Node Types | Search/Traversal Strategy | ||||||
| Contain | Import | Inherit | Invoke | Directory | File | Class | Function | ||
| CodexGraphLiu et al. (2024) | ✔ ✗ | ✗ | ✔ | ✔ | ✗ | ✗ | ✔ | ✔ | Cypher queries |
| RepoGraphOuyang et al. (2025) | ✔ ✗ | ✗ | ✔ | ✔ | ✗ | ✗ | ✔ | ✔ | Ego-graph retrieval |
| RepoUnderstanderMa et al. (2024) | ✔ | ✗ | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | MCTS |
| OrcaLocaYu et al. (2025) | ✔ | ✗ | ✗ | ✔ ✗ | ✔ | ✔ | ✔ | ✔ | Simple search tools |
| LocAgent(Ours) | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | Unified retrieval tools |
Due to the inherent structure of code, several works have employed graph-based representations to improve code understanding by capturing key relationships between components. Aider (2023) constructs a RepoMap and uses a graph ranking algorithm to identify the most significant contextual elements. Similarly, as a plugin, RepoGraph Ouyang et al. (2025) performs subgraph retrieval – extracting an ego-network of relevant lines and their neighbors – to provide structured context. CodexGraph Liu et al. (2024) indexes the repository into a Neo4j graph database, where LLM agents query the database precisely using Cypher. The efficiency of its retrieval process depends heavily on the querying capabilities of the LLM. These methods focus primarily on providing relevant context but do not enhance the traversal process itself, as they do not explicitly model directory structure or file hierarchies.
In contrast, RepoUnderstander Ma et al. (2024) builds hierarchical and function-call graphs, using Monte Carlo Tree Search (MCTS) guided by an LLM for exploration. While thorough, MCTS introduces extra computational overhead, making it less efficient than simpler traversal methods like BFS, particularly in large repositories. OrcaLoca Yu et al. (2025) uses a simplified graph enhanced by priority scheduling and context pruning. It maintains efficient search but may miss complex invocation dependencies. Table 1 summarizes the differences between these methods and LocAgent. Compared to these approaches, LocAgent offers a more comprehensive and unified representation of the repository, along with efficient, unified retrieval tools specifically designed for LLM consumption.
@@ 3.2 Agent-guided Code Search
3.3 Open-source Model Fine-tuning
@@ 4.2 Dataset Construction
| Loc-Bench | Bug Report | 242 |
| Feature Request | 150 | |
| (Totoal = 560) | Security Issue | 29 |
| Performance Issue | 139 |
@@ 5.5 Efficiency Analysis
Table 5 presents an efficiency analysis comparing agent-based methods in terms of cost and the number of agent interactions required. MoatlessTools demonstrates good cost-efficiency and requires relatively fewer rounds of interaction. However, the dense embeddings it uses make it difficult and slow to adapt to fast-evolving codebases. SWE-agent and Openhands also show moderate costs but still do not match the efficiency of LocAgent. For LocAgent with Claude-3.5, although more rounds of interaction are required, the cost remains lower than that of Openhands, illustrating the token efficiency of our tool’s outputs. LocAgent with fine-tuned Qwen models stands out for its superior efficiency111We calculate the cost based on the prices from AI inference providers Hyperbolic (2025); artificialanalysis.ai (2025). Specifically, for the Qwen2.5-32B(ft) model, the cost is $0.20/1M tokens for both input and output. For the Qwen2.5-7B(ft) model, the cost is $0.14/1M tokens for input and $0.28/1M tokens for output.. Qwen2.5-7B(ft) is the most cost-efficient option, requiring only $0.05 per example, while Qwen2.5-32B(ft) offers a more cost-effective alternative to Claude-3.5. These results highlight the potential of fine-tuned open-source models as efficient alternatives, providing an optimal balance of cost-effectiveness and performance that surpasses other methods.
@@ 5.6 Ablation Study
5.7 Evaluation Results on Loc-Bench
@@
| IR-Based | CodeRankEmbed | 74.29 | 80.89 | 63.21 | 67.50 | 43.39 | 46.61 |
| Agentless | Claude-3.5 | 67.50 | 67.50 | 53.39 | 53.39 | 42.68 | 42.68 |
| OpenHands | Claude-3.5 | 79.82 | 80.00 | 68.93 | 69.11 | 59.11 | 59.29 |
| SWE-agent | Claude-3.5 | 77.68 | 77.68 | 63.57 | 63.75 | 51.96 | 51.96 |
| LocAgent (Ours) | Qwen2.5-7B(ft) | 78.57 | 79.64 | 63.04 | 63.04 | 51.43 | 51.79 |
| Claude-3.5 | 83.39 | 86.07 | 70.89 | 71.07 | 59.29 | 60.71 | |
To ensure the robustness and generalization of our methods and fine-tuned Qwen models, and to eliminate potential data leakage, we evaluate our new dataset. Since Loc-Bench includes examples that edit 1 to 5 files, we assess file localization at top-5 and top-10 ranks, and function/module localization at top-10 and top-15 ranks. Table 7 shows that our fine-tuned Qwen2.5-7B model exhibits strong generalization capabilities, maintaining competitive performance compared to SWE-agent using more expensive and strong model. These results highlight the practicality of the fine-tuned Qwen2.5-7B model for real-world applications. Despite being an open-source alternative, it achieves a performance comparable to Claude-3.5, supporting its feasibility as a cost-effective substitute for commercial models in practical scenarios.
Additionally, we evaluate the performance across four different difficulty categories. Figure 5 clearly shows that our method outperforms other methods in almost all categories of code localization. However, it also highlights a noticeable decrease in performance across the other three categories compared to the Bug Report category. This performance gap likely reflects our training data distribution, which contained more bug report examples, potentially leading to scaffolds better optimized for bug localization tasks. This trend suggests that while our method is highly effective for bug report localization, there is still room for improvement in handling the other categories through more balanced training data and category-specific optimization strategies.